Examine the following solution to 𝑥 2 − 2𝑥 = −1: 𝑥(𝑥 − 2) = −1 𝑥 = −1 or 𝑥 − 2 = −1 𝑥 = −1 or 𝑥 = 1 Is this method correct? Explain. 


1
Expert's answer
2021-11-09T16:53:23-0500

QUESTION


Examine the following solution to x²− 2x = −1:

x(x − 2) = −1

x = −1

or x − 2 = −1

x = −1 or x = 1

is this method correct? explain.


ANSWER


The method is not correct

The root of the quadratic equation is "1" because it is a perfect square.


The above method have given two distinct values:

"-1 \> and \> 1"  and are not both the roots of quadratic function "x\u00b2-2x+1" 


When "x=-1" is substituted in the quadratic equation

"(-1)²-2(-1)+1\ne0"


The method is assumed to be same as "x(x-2)=0. \>"

In these case either "x=0\>or\>x-2=0"

This will always be true because:


For "\alpha\>and \>\beta\>\in\>\Reals"

when either is zero, then

"\alpha\beta=O"



But for  "x(x-2)=-1" and "x=-1"the equation can only be true if 

"x-2=1."

Otherwise this method will give invalid roots.

The method also treat the function

"f(x) =x(x-2) \>"

as a function with two variables.

In that for "for \>f(x)=-1"

the first "x=-1" and "(x-2)=1"

The second "x" in the expression becomes "3" so that "x(x-2)=-1"

The method is invalid